# EquiContact: A Hierarchical SE(3) Vision-to-Force Equivariant Policy for Spatially Generalizable Contact-rich Tasks

### v2

@@ Abstract
This paper presents a framework for learning vision-based robotic policies for contact-rich manipulation tasks that generalize spatially across task configurations. We focus on achieving robust spatial generalization of the policy for the peg-in-hole (PiH) task trained from a small number of demonstrations. We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant Action Chunking Transformers, G-CompACT). G-CompACT operates using only localized observations (geometrically consistent error vectors (GCEV), force-torque readings, and wrist-mounted RGB images) and produces actions defined in the end-effector frame. Through these design choices, we show that the entire EquiContact pipeline is $SE(3)$-equivariant, from perception to force control. We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance. Real-world experiments on PiH, screwing, and surface wiping tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles. The experimental videos can be found on the project website: https://sites.google.com/berkeley.edu/equicontact
@@ I Introduction
An alternative line of recent research focuses on leveraging symmetry—particularly equivariance—to enhance spatial generalizability, thereby improving sample efficiency during training [10, 11]. This approach requires fewer data but comes with its own challenges. Equivariant neural networks, being of a more specialized nature, are often not as well-developed and are more computationally intensive than their non-equivariant counterparts, making real-time and large-scale deployment more difficult. As a result, it becomes more attractive for users to use standard models trained with massive datasets in many instances.
In [11], a $SE(3)$-equivariant gain-scheduling policy using geometric impedance control (GIC) [12, 13] was proposed to solve peg-in-hole (PiH) problems. Inspired by the view that many manipulation tasks can be framed as pick-and-place problems [4], we modeled PiH as a compliant pick-and-place task, where final peg poses are provided by vision-based $SE(3)$-equivariant models such as Diffusion-EDF (Diff-EDF) [14]. Since both the high-level planner and low-level variable impedance controller are equivariant, they can be combined to form a vision-to-force equivariant policy. However, in practice, Diff-EDF’s placement accuracy proved insufficient for tight insertion tasks, which require sub-millimeter precision (details in Appendix -A). This revealed a key limitation: high-level vision planners may capture global structures but struggle with precision and contact-sensitive execution. From this insight, we introduce an intermediate layer between the planner and the low-level controller, which provides real-time visual feedback to correct the residual errors of the high-level planner.
@@
    We identify three key principles for spatially generalizable contact-rich manipulation: left-invariant compliant control action (via GAC [11]), localized policy (left invariance), and induced equivariance. These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [10].
@@
From these key principles, we propose a general framework to enhance the spatial generalization and interpretability of vision-based policies, namely, “anchoring localized policy on globally estimated reference frame.” We emphasize that our work provides complementary insights to recent trends of robot learning [1, 8, 16, 2] that aim to build generalist policies from a large-scale demo dataset. Our principles offer structural guidelines to improve spatial generalizability through $SE(3)$ equivariance.
@@ II Related Works
Visuomotor Servoing Methods Recently, generative modeling has become mainstream in realizing visuomotor servoing policies. Particularly, there are two dominant methods for visuomotor servoing: Action Chunking with Transformers (ACT) [6] and Diffusion Policy (DP) [5]. ACT utilizes the conditional variational autoencoder (CVAE) method as a generative model, whereas DP employs denoising diffusion as its generative model. ACT and DP have been extended to other approaches, including compliance and force-reactive behaviors [17, 18, 19], as well as structural improvements [20, 16, 21]. Our work is most closely related to CompliantACT (CompACT) [17], which integrates compliant control for visuomotor policies. We have significantly improved CompACT by incorporating a provable $SE(3)$ equivariant structure.
@@
Figure 1: We propose an EquiContact, a hierarchical, provably $SE(3)$ vision-to-force equivariant policy for spatially generalizable contact-rich tasks. The proposed EquiContact consists of G-CompACT (Section. IV-B) and Diffusion-EDF (Section. IV-A). The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen task transformations during evaluation. The overall EquiContact algorithm is summarized in Algorithm 1.
@@ III Problem Definition
Unlike prior work [11] that assumes a known hole pose and a pre-grasped peg, we consider a more general setup: the robot must first grasp the peg and then perform insertion using only vision and proprioception, as illustrated in Fig. 2. We assume only that the peg is upright in the scene. Inspired by the idea that many manipulation tasks can be viewed as sequences of pick-and-place actions, we formulate insertion as a compliant pick-and-place task. However, high-level vision planners often lack the precision needed for tight-tolerance tasks like PiH, which in our case requires $1\mathrm{mm}$ accuracy.
@@
To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location, where its objective is to train a policy that performs nearly perfectly in the trained scenario. We then evaluate benchmark and proposed methods—trained solely on these limited demonstrations—across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented. 22 external cameras with calibrated extrinsics and 22 wrists cameras are installed. (Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance. (Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
@@ IV Solution Approach
We now describe our proposed EquiContact policy, which integrates a high-level vision planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant ACT (G-CompACT). G-CompACT consists of a behavior cloning module based on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC). As discussed in Section III, we focus on the insertion task, with extension to picking addressed later in the paper. The Diff-EDF and G-CompACT are trained separately and combined at inference, as shown in Fig. 1.
@@
Conceptually, EquiContact follows a simple yet powerful principle: “anchoring a localized policy on a globally estimated reference frame.” In our framework, G-CompACT serves as the localized low-level policy, operating solely on observations defined in the current end-effector frame; wrist camera RGB images, force-torque (FT) signals, and a geometrically consistent error vector (GCEV) $e_{{}_{G}}$. It outputs the desired relative pose and admittance gains, also in the end-effector frame. As both inputs and outputs are defined locally, G-CompACT functions as a fully localized policy. The high-level planner, Diff-EDF, estimates the pose of the target (e.g., hole) in the world frame. At the inference, the robot moves near the estimated reference frame, and G-CompACT is activated to perform compliant motion using only local feedback. Because the low-level policy does not depend on absolute global inputs, it can transfer robustly to unseen spatial configurations when the estimated reference frame is provided. This divide-and-conquer design provides a general framework for enhancing both spatial generalization and policy interpretability for contact-rich, and more broadly, general manipulation tasks.
In the remainder of this section, we formalize this spatial generalization property as $SE(3)$ equivariance and show how EquiContact satisfies this by construction through a rigorous mathematical derivation. We note that $\theta$, $\theta_{i}$, and $\phi$ with $i=1,2,3$ denote a learnable parameter of the neural network, without loss of generality. We also note that we follow the notation and definition conventions of [11, 30].
IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
Diff-EDF [14] is a diffusion-based manipulation policy with a bi-equivariant structure on $SE(3)$ transformations for pick-and-place tasks. Diff-EDF takes two point clouds as input: the scene point cloud $\mathcal{O}^{scene}$ captured by two external RGBD cameras with calibrated extrinsics (Fig. 2), and the gripper point cloud $\mathcal{O}^{grasp}$ in the end-effector frame, with or without a grasped object.
Diff-EDF outputs a pose $g_{{}_{EDF}}\in SE(3)$ representing the estimated target pose of the object of interest, which in our case is the hole position and is defined as:
$g_{{}_{EDF}}=f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$ (1)
where the neural network $f_{\theta_{1}}$ implements the mapping. Due to the left-equivariant property, it follows that: $g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
Let $\mathcal{O}^{ref}\subset\mathcal{O}^{scene}$ denote the subset of points corresponding to the object of interest (e.g., the hole structure, see also Fig. 3). The model is designed to satisfy the following left-equivariance property on the point cloud characterization of the target object of interest:
Assumption 1 (Left-Equivariance of Diff-EDF).
$\begin{split}g_{{}_{EDF}}&=f_{\theta_{1}}(\mathcal{O}^{ref},\mathcal{O}^{grasp})\\ &\implies g_{l}\cdot g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref},\mathcal{O}^{grasp}),\end{split}$ (2)
To meet this assumption, we randomize the pose of the hole-platform assembly during training, both translationally and rotationally, with visual distractors in the background. This encourages the model to focus on $\mathcal{O}^{ref}$, not the rest of the objects in the scene. While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and has high robustness to left $SE(3)$ transformations.
To train Diff-EDF, the scene and grasp point clouds are collected together with the target reference frames, which represent the desired poses of the end-effector for pick-and-place operations. $15$ demonstrations were collected for the Diff-EDF: $12$ samples of the flat platform and $3$ samples of the tilted platform, both translationally and rotationally randomized. The training process of Diff-EDF follows the procedure in [14].
Note that arbitrary peg transformations relative to the gripper can be handled via the right-equivariance feature of the Diff-EDF; however, due to the complications mentioned in Section. III, we enforced a consistent grasp and only utilized the left-invariance property. We refer the readers to [14, 10] for full details of Diff-EDF.
@@ IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
IV-B Geometric Compliant control Action Chunking with Transformers (G-CompACT)
IV-B1 G-CompACT
@@
G-CompACT is based on the Action Chunking with Transformer (ACT), which is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks [6]. To make G-CompACT spatially equivariant, we follow the principles proposed in [11]: left-invariant policy and policy representation in the end-effector body frame. To achieve this, we structure the observations and actions of the G-CompACT as described below.
@@
First, the observations are given by, (1) Geometrically Consistent Error Vector (GCEV) $e_{{}_{G}}$ proposed in [11], (2) FT sensor in the end-effector frame $F_{e}$ to capture contact behaviors, and (3) RGB images $I_{w}=\{I_{w,1},I_{w,2}\}$ from wrist cameras (see Fig. 3). For actions, we choose (1) relative pose from the current end-effector frame $g_{rel}$, and (2) admittance gains for Geometric Admittance Control (GAC) $(K_{p},K_{R})$. The details of GAC and the definition of gains will be provided later in the Section.
Formally, let the G-CompACT method be $\pi_{\theta_{2}}$ such that
$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq(g_{rel},K_{p},K_{R})(k),\quad o(k)\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (3)
@@
where $a(k)$ denotes the actions, and $o(k)$ denotes the observation at time step $k$. In fact, our G-CompACT also outputs the actions of chunk size $N$ following the standard action-chunking strategy. However, for the compact presentation, we will only consider the single-step action after undergoing proper processing, such as a temporal ensemble. The GCEV $e_{{}_{G}}(g,g_{{}_{EDF}})$ is defined as
$e_{{}_{G}}(g,g_{{}_{EDF}})=\begin{bmatrix}R^{T}(p-p_{{}_{EDF}})\\ (R_{{}_{EDF}}^{T}R-R^{T}R_{{}_{EDF}})^{\vee}\end{bmatrix}$, (4)
@@
where $g=(p,R)\in SE(3)$ is a current end-effector pose, $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF, and $(\cdot)^{\vee}$ denotes the vee-map, a mapping from $so(3)$ (Lie algebra of $SO(3)$) to $\mathbb{R}^{3}$. The physical meaning of GCEV is an error vector in the current end-effector frame between the current end-effector pose and the reference frame, considering the $SE(3)$ manifold structure, i.e., non-Euclidean. For the details of GCEV $e_{{}_{G}}$, we refer to [11, 12].
@@
Since the images $I_{w}$ are fed to the transformer decoder after being passed to the visual encoder structure, one can further consider the G-CompACT as
$\begin{split}a(k)&=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)(k),\quad\text{where}\;z(k)=\mu_{\phi}(I_{w}).\end{split}$ (5)
where $\mu_{\phi}$ denotes the visual encoder, e.g., ResNet, and $z$ is the latent variable from the visual encoder.
We now introduce the following assumption regarding the left-invariant visual representation.
Assumption 2 (Left-invariant Visual Features).
@@
The features from visual encoder $\mu_{\phi}$ are left invariant, i.e.,
$z(k)=\mu_{\phi}(g_{l}\circ I_{w})=\mu_{\phi}(I_{w}),\;\;\forall g_{l}\in SE(3).$ (6)
Note that we refer to $g_{l}$ as a left-group action [10]. The left-group action applied on the wrist-camera images is illustrated in Fig. 3. The meaning of the visual representation $z$ being left-invariant is that the vision encoder $\mu_{\phi}$ is trained to focus only on group action invariant features, such as the flat surface surrounding the hole on the platform. The satisfaction of Assumption 2 can be challenging. To satisfy this assumption, we have designed our platform and the surface surrounding the hole assembly so that there are sufficient surface features and the cameras primarily see the surface surrounding the hole, not the lower parts of the platform. This engineering choice was somewhat ad hoc, as there is no guarantee or sufficient inductive bias to encourage the desired behavior. We will demonstrate in later experimental results that this assumption may not apply in certain cases.
@@
Figure 3: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$. As the left group action is applied on the end-effector, the wrist cameras start to see backgrounds other than the optical table. $\{s\}$ denotes a spatial frame, i.e., robot base frame or world frame.
@@
The following proposition shows the left-invariance of the G-CompACT method in the end-effector frame.
@@
Proposition 1 (Left-invariance of G-CompACT).
Suppose that the Assumption 2 holds. Then,
$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi_{\theta_{2}}(o(k)).$ (7)
@@
The proof is presented in Appendix. -B. In what follows, we present an equivariant property of the pose signal produced by G-CompACT when described in the spatial frame.
@@
The G-CompACT $\pi_{\theta_{2}}$ represented in the spatial frame satisfies the following equivariance property:
$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\pi_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (8)
Proof.
Notice that the desired pose signal in the spatial frame $g_{d}$ is obtained via $g_{d}=gg_{rel}=g\pi_{\theta_{2}}(o(k))$. Then, utilizing the left-invariance property, $g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ o(k))=g_{l}g_{d}$. ∎
To train a G-CompACT, we collect the expert demonstration using teleoperation on the fixed pose of the platform. During data collection, the expert teleoperator monitors the task’s progress, makes real-time movement commands via a SpaceMouse, and adjusts the admittance gains using keyboard input to switch between the following predefined gain modes: low-gain mode, high-gain mode, insertion mode, and contact mode. The low/high gain mode has low/high gains in all directions, the insertion mode has high gains in the $z$ direction of the end-effector frame and low gains elsewhere. Finally, the contact mode has low gains in the $z$ direction and high gains elsewhere. We collected $86$ demonstrations to train a near-perfect policy.
Since we know the fixed pose of the platform a priori, e.g., a ground-truth reference frame, the GCEV vector can be calculated for the training process. Nevertheless, the reference frame needs to be estimated via Diff-EDF (as $g_{{}_{EDF}}$) during the inference stage, which may have non-negligible errors. To handle this issue, we have added noise to the reference frame $g_{ref}$ to calculate $e_{{}_{G}}$ during dataset preprocessing. This provides the model with an inductive bias to primarily rely on $e_{{}_{G}}$ values for rough alignment and rely on vision feedback for fine-grained motion. The rest of the training follows the standard imitation learning pipeline.
IV-B2 Geometric Admittance Control (GAC)

### v3

@@ Abstract
This paper presents a framework for learning vision-based robotic policies for contact-rich manipulation tasks that generalize spatially across task configurations. We focus on achieving robust spatial generalization of the policy for the contact-rich tasks trained from a small number of demonstrations. We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant Action Chunking Transformers, G-CompACT). G-CompACT operates using only localized observations (geometrically consistent error vectors (GCEV), force-torque readings, and wrist-mounted RGB images) and produces actions defined in the end-effector frame. Through these design choices, we show that the entire EquiContact pipeline is $SE(3)$-equivariant, from perception to force control. We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance. Real-world experiments on peg-in-hole (PiH), screwing, and surface wiping tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles. The experimental videos will be attached as supplementary material, and the codes will be released.
@@ I Introduction
An alternative line of recent research focuses on leveraging symmetry—particularly equivariance—to enhance spatial generalizability, thereby improving sample efficiency during training [22, 19]. This approach requires less data but comes with its own challenges. Equivariant neural networks, being of a more specialized nature, are often not as well-developed and are more computationally intensive than their non-equivariant counterparts, making real-time and large-scale deployment more difficult. As a result, it becomes more attractive for users to use standard models trained with massive datasets in many instances.
In [19], a $SE(3)$-equivariant gain-scheduling policy using geometric impedance control (GIC) [20, 21] was proposed to solve peg-in-hole (PiH) problems. Inspired by the view that many manipulation tasks can be framed as pick-and-place problems [24], we modeled PiH as a compliant pick-and-place task, where final peg poses are provided by vision-based $SE(3)$-equivariant models such as Diffusion-EDF (Diff-EDF) [18]. Since both the high-level planner and low-level variable impedance controller are equivariant, they can be combined to form a vision-to-force equivariant policy. However, in practice, Diff-EDF’s placement accuracy proved insufficient for precision tasks, which require sub-millimeter precision (details provided in Appendix AIII-A. This revealed a key limitation: high-level vision planners may capture global structures but struggle with precision and contact-sensitive execution. Henceforth, we introduce an intermediate layer between the planner and the low-level controller, which provides real-time visual feedback to correct the residual errors of the high-level planner.
@@
    We identify three key principles for spatially generalizable contact-rich manipulation: (1) left-invariant compliant control action (via GAC [19]), (2) localized policy (left invariance), and (3) induced equivariance. These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [22].
@@
From these key principles, we propose a general framework to enhance the spatial generalization and interpretability of vision-based policies, namely, “anchoring localized policy on globally estimated reference frame.” We emphasize that our work provides complementary insights to recent trends in robot learning [13, 9, 5, 1] that aim to build generalist policies from large-scale demonstration datasets. Our principles provide structural guidelines for improving spatial generalizability via $SE(3)$ equivariance.
@@ II Related Works
Visuomotor Servoing Methods Recently, generative modeling has become mainstream in realizing visuomotor servoing policies. Particularly, there are two dominant methods for visuomotor servoing: Action Chunking with Transformers (ACT) [37] and Diffusion Policy (DP) [4]. ACT uses a conditional variational autoencoder (CVAE) as its generative model, whereas DP uses denoising diffusion. ACT and DP have been extended to other approaches, including compliance and force-reactive behaviors [8, 7, 32], as well as structural improvements [14, 5, 29]. Our work is most closely related to CompliantACT (CompACT) [8], which integrates compliant control for visuomotor policies. We have significantly improved CompACT by incorporating a provable $SE(3)$ equivariant structure.
@@ III Problem Definition
Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented. 22 external cameras with calibrated extrinsics and 22 wrists cameras are installed. The workspace shown is the Diff-EDF workspace. (Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance. (Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
@@
Unlike prior work [19] that assumes a known hole pose and a pre-grasped peg, we consider a more general setup: the robot must first grasp the peg and then perform insertion using vision, proprioception, and task description in text, as illustrated in Fig. 2. We assume the peg is upright and that the hole’s yaw angle is known within $90^{\circ}$ range. Given an initial estimate, we resolve the orientation by selecting the closest angle among the four symmetric candidates (e.g., $\psi,\psi+90^{\circ},\psi+180^{\circ},\psi+270^{\circ}$). As mentioned earlier, high-level vision planners often lack the precision needed for tight-tolerance tasks like PiH, which in our case requires $1\mathrm{mm}$ accuracy.
@@
To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location, where its objective is to train a policy that performs nearly perfectly in the trained scenario. We then evaluate benchmark and proposed methods, trained solely on these limited demonstrations, across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy. Further, we demonstrate that our proposed approach can adapt to extreme task transformations as shown in Fig. 3.
Finally, the validity of the proposed EquiContact method is shown in other contact-rich tasks, such as the screwing task and surface-wiping tasks.
Figure 3: Extreme task transformations. (Left) $90^{\circ}$ transformation in $x$ axis. (Middle) $135^{\circ}$ transformation in $x$ axis. (Right) $45^{\circ}$ transformation in $y$ axis, facing towards the camera.
@@ IV Solution Approach
We introduce the EquiContact framework in this Section. The EquiContact framework integrates a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) with a low-level compliant visuomotor policy (Geometric Compliant ACT, G-CompACT) and geometric admittance control (GAC) at the lowest level. The Diff-EDF gets the point cloud inputs from external cameras to generate reference frames. Based on the estimated reference frames, the G-CompACT process the real-time proprioceptive and wrist camera feedback to output desired poses and admittance gains. In what follows, the GAC module outputs the geometrically consistent compliant motion from desired poses and admittance gains to enable equivariant force interaction. We first focus on the insertion task, with extension to picking addressed later in the paper.
@@
Conceptually, EquiContact follows a simple yet powerful principle: “anchoring a localized policy on a globally estimated reference frame.” In our framework, G-CompACT serves as a fully localized low-level policy, operating solely on observations defined in the current end-effector frame, and actions defined in the end-effector frame – See Fig. 4. The high-level planner, Diff-EDF, estimates the pose of the target (e.g., hole) in the world (global) frame. At the inference, the robot moves near the estimated reference frame, and G-CompACT is activated to perform compliant motion using only local feedback. Because the low-level policy does not depend on absolute global inputs, it can transfer robustly to unseen spatial configurations when the estimated reference frame is provided. This divide-and-conquer design provides a general framework for enhancing both spatial generalization and policy interpretability for contact-rich, and more broadly, general manipulation tasks. In the remainder of this section, we formalize this spatial generalization property as $SE(3)$ equivariance and show how EquiContact satisfies this by its design choices.
@@
IV-A Geometric Compliant control Action Chunking with Transformers (G-CompACT)
@@
G-CompACT is based on the Action Chunking with Transformer (ACT), which is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks [37]. To make G-CompACT spatially equivariant, we follow the principles proposed in [19]:
  • Left-invariant policy, and
  • Policy representation in the end-effector body frame
We have designed the G-CompACT architecture to achieve this properties as described in this chapter. The overall structure of the G-CompACT is also summarized in Fig. 4.
Figure 4: G-CompACT architecture is presented. The G-CompACT $\pi_{\theta}$ receives GCEV $e_{{}_{G}}$ (2) and F/T sensor value $F_{e}$ as proprioceptive inputs, along with two wrist camera images $I_{w}$. The wrist camera images are fed to the CLIP-ResNet50 visual backbone, followed by the Feature-wise Linear Modulation (FiLM) layer. The FiLM layer is modulated by the text tokens $t_{text}$ from the (frozen) CLIP text encoder, which processes the task descriptions. The proprioceptive inputs and the latent features of modulated vision $z$ are then fed to the transformer decoder $D_{\psi}$, which outputs the action signals $a$. Note that we omitted the style variable for the transformer decoder, and 0 values are used during inference.
@@
The observations are given by, (1) Geometrically Consistent Error Vector (GCEV) $e_{{}_{G}}$ proposed in [19], (2) FT sensor in the end-effector frame $F_{e}$ to capture contact behaviors, and (3) RGB images $I_{w}=\{I_{w,1},I_{w,2}\}$ from wrist cameras (see Fig. 5). For actions, we choose (1) relative pose from the current end-effector frame $g_{rel}$, and (2) admittance gains for Geometric Admittance Control (GAC) $(K_{p},K_{R})$. The details of GAC and the definition of gains will be provided later in the Section. Formally, the G-CompACT method $\pi_{\theta}$ can be written as:
$\begin{split}a(k)&=\pi_{\theta}(o(k)),\quad\text{where}\\ a(k)&\triangleq(g_{rel},K_{p},K_{R})(k),\quad o(k)\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (1)
@@
where $a(k)$ denotes the actions, and $o(k)$ denotes the observation at time step $k$. Although the G-CompACT outputs the actions of chunk size $N$, we will only consider the single-step action after proper processing, such as a temporal ensemble, for notational compactness. The GCEV $e_{{}_{G}}(g,g_{ref})\in\mathbb{R}^{6}$ is defined as
$e_{{}_{G}}(g,g_{ref})=\begin{bmatrix}R^{T}(p-p_{ref})\\ (R_{ref}^{T}R-R^{T}R_{ref})^{\vee}\end{bmatrix}$, (2)
@@
where $g=(p,R)\in SE(3)$ is a current end-effector pose, $g_{ref}=(p_{ref},R_{ref})\in SE(3)$ is the reference frame estimated by the global estimator, e.g., Diff-EDF, and $(\cdot)^{\vee}$ denotes the vee-map, a mapping from $so(3)$ (Lie algebra of $SO(3)$) to $\mathbb{R}^{3}$. The physical meaning of GCEV is an error vector between the current end-effector frame and the reference frame, defined on the current end-effector frame. As will be elaborated in Appendix AI, the proprioceptive signals $e_{{}_{G}}$ and $F_{e}$ are left-invariant. For the details of GCEV $e_{{}_{G}}$, we refer to [19, 20].
@@
The images $I_{w}$ are fed to the transformer decoder $\mathcal{D}_{\psi}$ after being processed by the vision encoder structure $\mu_{\phi}$; therefore, one can further represent G-CompACT as
$a(k)=\mathcal{D}_{\psi}(e_{{}_{G}},F_{e},z)(k)=\mathcal{D}_{\psi}(e_{{}_{G}},F_{e},\mu_{\phi}(I_{w}))(k),$ (3)
where $z=\mu_{\phi}(I_{w})$ is a visual feature from the vision encoder. To satisfy the left-invariant condition of G-CompACT, the features from the vision encoder $z$ need to be invariant to the left-transformation of the image. We formalize the left-invariant visual feature condition as the following assumption.
Assumption 1 (Approximately Left-invariant Visual Features).
@@
The visual encoder $\mu_{\phi}$ produces features that are approximately left-invariant to task transformations, i.e.,
$\mu_{\phi}(g_{l}\circ I_{w})\approx\mu_{\phi}(I_{w}),$ (4)
$\forall g_{l}\in SE(3)$ that preserves local task geometry.
Here, $\approx$ notation denotes invariance up to a bounded representation error that does not affect the policy’s qualitative behavior. Note that we refer to $\circ$ as a group action [22]. The left-group action applied to the wrist-camera images $g_{l}\circ I_{w}$ is illustrated in Fig. 5. The meaning of the visual representation $z$ being left-invariant is that the vision encoder $\mu_{\phi}$ is trained to focus only on group action invariant features, such as the flat surface surrounding the hole on the platform. To satisfy this assumption, we use language grounding to extract vision features that are correlated with the language description. The core insight behind this is that language tokens encode object identity rather than pose, and thus provide a conditioning signal that is invariant to global $SE(3)$ transformations of the scene. For example, a “peg” is still “peg” no matter from which view it is seen.
Specifically, the pretrained CLIP-ResNet5050 (CLIP-RN50) [16] is utilized for the vision encoder, and the CLIP text encoder is also employed. Although the pretrained CLIP-RN50 was used, it was fully retrained (details provided in Appendix AII); thereby, it served as a good initialization. We provide the task descriptions to the text encoder to obtain text tokens $t_{text}$. The vision feature $z_{raw}$ of CLIP-RN50 backbone is then modulated using feature-wise linear modulation (FiLM, [15]) layer from $t_{text}$ via
$z=\beta(t_{text})z_{raw}+\gamma(t_{text}),$ (5)
where $\beta$ and $\gamma$ are trainable FiLM layer. Using FiLM, the vision features are suppressed or highlighted to align with task-relevant semantic concepts, empirically encouraging approximate left-invariance with respect to workspace transformations.
@@
Figure 5: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$. As the left group action is applied to the end-effector and the target object, the wrist cameras start to see the backgrounds of arbitrary lab objects.
@@
Under the satisfaction of Assumption 1, i.e., ideal left-invariant vision feature condition, the following proposition shows the left-invariance of the G-CompACT method in the end-effector frame.
@@
Proposition 1 (Left-invariance of G-CompACT under ideal invariant visual features).
Suppose Assumption 1 holds. Then the G-CompACT policy $\pi_{\theta}$ is left-invariant in the end-effector frame, i.e.,
$\pi_{\theta}(g_{l}\circ o(k))=\pi_{\theta}(o(k)),\quad\forall g_{l}\in SE(3).$ (6)
The proof is presented in Appendix AI. A remark is provided.
Remark 1 (Implication of Proposition 1).
Proposition 1 highlights a sufficient structural condition for spatial generalization: if the visual encoder and proprioceptive signals can be made (approximately) left-invariant to $SE(3)$ task transformations, then the resulting closed-loop policy inherits left-invariance by construction. This motivates learning or enforcing representations that satisfy Assumption 1.
@@
In what follows, we present an equivariant property of the pose signal produced by G-CompACT when described in the spatial frame.
@@
The G-CompACT $\pi_{\theta}$ represented in the spatial frame satisfies the following equivariance property:
$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\pi_{\theta}(g_{l}\circ o(k)).\end{split}$ (7)
The proof is presented in Appendix AI.
IV-B Geometric Admittance Control (GAC)